### v1

# Denoising Nearest Neighbor Graph via Continuous CRF for Visual Re-ranking without Fine-tuning

Abstract
Visual re-ranking using Nearest Neighbor graph (NN graph) has been adapted to yield high retrieval accuracy, since it is beneficial to exploring an high-dimensional manifold and applicable without additional fine-tuning. The quality of visual re-ranking using NN graph, however, is limited to that of connectivity, i.e., edges of the NN graph. Some edges can be misconnected with negative images. This is known as a noisy edge problem, resulting in a degradation of the retrieval quality. To address this, we propose a complementary denoising method based on Continuous Conditional Random Field (C-CRF) that uses a statistical distance of our similarity-based distribution. This method employs the concept of cliques to make the process computationally feasible. We demonstrate the complementarity of our method through its application to three visual re-ranking methods, observing quality boosts in landmark retrieval and person re-identification (re-ID).
1 Introduction
(a) Initial NN graph
(b) NN graph after reciprocity check
(c) Diffusion-based retrieval of (b)
(d) NN graph denoised by ours
(e) Diffusion-based retrieval of (d)
Figure 1: This figure shows NN graph (affinity) of toy data on manifold. The blue and green colors of nodes indicate different manifolds, while a yellow node and red nodes represent a query point and the retrieval results from a visual re-ranking method of diffusion. The red edges are noisy edges that are connected between different manifolds. Results (e) of our approach (d) are on the correct manifold. Best viewed with zoom-in.
Nearest neighbor search approaches  [1, 2, 3] have used various visual re-ranking techniques  [4, 5, 6] for further improvement of retrieval accuracy. Among them, average query expansion (AQE)  [4] was considered to explore an image manifold on the descriptor space in a simple manner by using a new query averaged over nearest neighbor features at a query time. However, AQE is limited to only information between query and database images.
Contrary to AQE, NN graph based retrieval methods, e.g., diffusion process [7, 8, 6, 9] and graph traversal [10, 11], fully makes use of an affinity matrix between database images as well as information between the query and database images. Diffusion methods, revisited from Donoser et al. [8], constructed the affinity matrix, while considering the query as a part of database images, and calculated solutions of diffusion methods in an iterative manner. However, regarding the query as one of database images requires huge computational costs in online time, because of creating a new affinity matrix whenever new queries are given. Iscen et al. [6] proposed a method that does not have to recreate the matrix even if new queries are given. Yang et al. [12] also made the Iscen’s method more efficient by shifting the diffusion step from online computing to offline one. Recently, Ouyang et al. [13] utilizes self-attention modules to aggregate similarities for improving visual re-ranking outcomes, while requiring additional fine-tuning. Shao et al. [14] highlight visual re-ranking from only top-M results using rich global features to enhance only retrieval precision, not recall. The reliance of this method on the initial ranked list may reduce its robustness across varied or broader retrieval contexts. Our work assumes conditions without fine-tuning and without rich global features, performing re-ranking across the entire database [15].
Yang et al. [16] tackled an issue where NN graph based re-ranking is quite sensitive to noisy data, since all edges, even those containing noise, are used to represent the affinity matrix. Thus, a locally constrained affinity construction was proposed based on edges of satisfying $k$ nearest neighbors ($k$-NN) constraints of images for relaxing noisy edges of negative images. A seminal approach [6] suggested an extension by filtering out noisy edges even from the locally constrained affinity by using a reciprocity check of outgoing and incoming edges of the locally constrained affinity (NN graph).
In this paper, we found that this reciprocity check with NN graph is insufficient to address the noisy edge problem, since the reciprocity check considers only reciprocal relationships between two edges. Since the reciprocity check and $k$-NN constraints do not fully exploit all information from each other, we address the problem with full consideration of relationships between all edges in a clique where we can conservatively reach a consensus (Fig. 1). To realize this, we efficiently employ a fully connected C-CRF, when denoising the noisy NN graph in offline time.
Conditional Random Field (CRF) has been extensively studied [17], and Continuous CRF (C-CRF) [18] was proposed for regression tasks such as depth estimation, image denoising and metric learning for person re-ID [19, 20, 21]. Departing to these applications, our approach targets to refine similarities between each of database images that are robust to noisy edges for improving a diffusion performance in offline time.
Main contributions are summarized as follows:
  • We introduce a C-CRF based refinement method for noisy NN graphs that serves as a pre-processing module to enhance visual re-ranking.
  • Our approach utilizes a statistical distance metric to ensure robust distance measurement and employs cliques for cost-effective denoising.
  • We show that our method significantly enhances existing visual re-ranking methods in landmark retrieval and person re-identification, confirming its role as a complementary tool.
2 Visual re-ranking with NN graph
In this section, we summarize the diffusion process to show how the NN-graph can be applied for visual re-ranking. The diffusion process in visual re-ranking constructs revised similarity considering manifold from initial similarity computed from all images in the database. Generally, this process constructs an undirected graph structure from the affinity matrix, which is defined using the pairwise similarity between images, and performs diffusion on this graph using a random walk formulation [6, 12].
We define a database $\mathcal{X}$ as a set of $N$ image descriptors, where each $\mathbf{x}_{i}$ denotes a descriptor of image $i$ with dimensionality $d$; $\mathbf{x}_{i}\in\mathbb{R}^{d}$. We define the affinity matrix $\mathbf{A}=(a_{ij})\in\mathbb{R}^{N\times N}$, $\forall i,j~{}\in~{}\{1,\dotsc,N\}$, following the reciprocity check with a local constraint by considering $k$ nearest neighbors [6], where each element is obtained by:
$s_{k}(\mathbf{x}_{i}|\mathbf{x}_{j})=\begin{cases}s(\mathbf{x}_{i},\mathbf{x}_{j})&\mathbf{x}_{i}\in\mathrm{NN}_{k}(\mathbf{x}_{j})\\ 0&\mathrm{otherwise}\end{cases},$ (1)
$a_{ij}=\min\{s_{k}(\mathbf{x}_{i}|\mathbf{x}_{j}),s_{k}(\mathbf{x}_{j}|\mathbf{x}_{i})\},$ (2)
where the similarity function $s(\cdot,\cdot)$ is positive and has zero self-similarity, and $\mathrm{NN}_{k}(\mathbf{x})$ denotes $k$-NN of $\mathbf{x}$. Intuitively, $a_{ij}$ equals to $s(\mathbf{x}_{i},\mathbf{x}_{j})$ if $\mathbf{x_{i}}$, $\mathbf{x_{j}}$ are the $k$ nearest neighbors of each other, and zero otherwise.
The degree matrix $\mathbf{D}$ is a diagonal matrix with the row-wise sum of $\mathbf{A}$. It is used to symmetrically normalize $\mathbf{A}$ as: $\mathbf{S}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$. After the matrix computation, starting from an arbitrary vector $\mathbf{v}^{0}$ that represents the initial similarity of a given query $\mathbf{y}$, the diffusion is performed until its state converges with a random walk iteration:
$\mathbf{v}^{t+1}=\rho\mathbf{Sv}^{t}+(1-\rho)\mathbf{v}^{0},~{}\rho\in(0,1).$ (3)
Zhou et al. [7] shows this iteration converges to a closed-form solution when assuming $10<\rho<$ 1:
$\mathbf{v}^{*}=(1-\rho)(\mathbf{I}-\rho\mathbf{S})^{-1}\mathbf{v}^{0}.$ (4)
In the conventional diffusion process, the values in $\mathbf{f}^{*}$ contain the refined similarity of the given query $\mathbf{y}$. Generally, the initial state $\mathbf{f}^{0}$ represents the similarities between the query $\mathbf{y}$ and the corresponding $k$ nearest neighbors as:
$\mathbf{v}^{0}=s_{k}(\mathbf{x}_{i}|\mathbf{y}),~{}~{}\forall i\in\{1,\dots,N\}.$ (5)
In this paper, we propose a novel, denoising approach for the initial affinity matrix to reduce the noisy edges.
3 Approach
Diffusion methods [6, 12] for image retrieval adopted a simple approach of using the reciprocity check (Eq. 2) with $k$-NN of each database image to ameliorate noisy edges whose vertices are not positive, i.e., similar images. We, however, found that the reciprocity check is limited to handling noise, since identifying the noisy edge is based on the consensus of only two nodes of an edge, resulting incorrect identification of noise edges.
While this noisy edge problem was not seriously treated in diffusion approaches for image retrieval, we tackle the noisy edge problem by considering all neighbors from each image in a well-connected subgraph where we can robustly detect noisy edges. Specifically, we realize our goal by using C-CRF on subgraphs, i.e., cliques, with our proposed weight function (illustrated in Fig. 2).
Figure 2: An example of clique-based denoising; $k=7$ and $L=10$. (a) shows initial edges of a yellow pivot node, and red edges are noisy edges. (b) shows a clique of the pivot, and all nodes in the clique are fully connected. (c) represents edges denoised by C-CRF, which is constructed from the clique.
3.1 C-CRF on clique
Our goal is to refine initial similarities of NN graph that are commonly set by top-$k$ similarities of each database image. For computational efficiency and effective noise handling, we utilize the concept of cliques and apply C-CRF into each of cliques. Specifically, we choose a clique, a complete subgraph, where we can have ample information such that we can identify noisy edges. On top of that, using the clique has computational benefits in running CRFs (Sec. 3.2).
We suppose that the whole database size is $N$ and a clique of a pivot image $\left\{{{I_{p}}}\right\}_{p=1}^{N}$ is ${{\cal C}_{p}}=\left({{{\cal N}_{p}},{{\cal E}_{p}}}\right)$, where a set of nodes (images) ${\cal N}_{p}$ is composed of $L$ nearest neighbors from the pivot image $I_{p}$ by a specified clique size $L$, and ${\cal E}_{p}$ is a set of edges that makes the clique ${\cal C}_{p}$, i.e., a complete subgraph. Given a clique ${\cal C}_{p}$, its set of initial similarities can be computed by the cosine similarity or Euclidean distance. For simplicity, we express $s(\mathbf{x}_{i},\mathbf{x}_{j})$ as ${s_{i,j}}$. The set is fed as input to C-CRF and is defined: ${{\bf{S}}_{{{\cal C}_{p}}}}=\left\{{{s_{i,j}}|i\in{{\cal N}_{p}},j\in{{\cal N}_{p}}}\right\}$ with the given initial similarity $s_{i,j}$ on clique ${\cal C}_{p}$. C-CRF computes a set of refined similarities defined as ${{\bf{y}}_{p}}=\left\{{{y_{p,i}}|i\in{{\cal N}_{p}}}\right\}$, which has $L$ elements with refined similarity $y_{p,i}$.
C-CRF on each clique ${\cal C}_{p}$ is represented by a conditional probability as follows:
$\begin{array}[]{l}{{\rm{P}}_{p}}\left({{{\bf{y}}_{p}}|{{\bf{S}}_{{{\cal C}_{p}}}}}\right)=\frac{1}{{Z\left({{{\bf{S}}_{{{\cal C}_{p}}}}}\right)}}\exp\left\{{-E\left({{{\bf{y}}_{p}},{{\bf{S}}_{{{\cal C}_{p}}}}}\right)}\right\},\end{array}$ (6)
[…]
$\begin{array}[]{l}\alpha\sum\nolimits_{i\in{{\cal N}_{p}}}{\phi\left({{y_{p,i}},{{\bf{S}}_{{{\cal C}_{p}}}}}\right)+\frac{\beta}{2}\sum\nolimits_{\left({i,j}\right)\in{{\cal E}_{p}}}{\psi\left({{y_{p,i}},{y_{p,j}},{{\bf{S}}_{{{\cal C}_{p}}}}}\right)}}.\end{array}$ (7)
The unary potential, $\phi\left({{y_{p,i}},{{\bf{S}}_{{{\cal C}_{p}}}}}\right)={\left({{y_{p,i}}-{s_{p,i}}}\right)^{2}}$, intends $y_{p,i}$ to follow the initial similarity $s_{p,i}$. To consider inter-image relationships, the pairwise potential is defined as:
$\psi\left({{y_{p,i}},{y_{p,j}},{{\bf{S}}_{{{\cal C}_{p}}}}}\right)=\textrm{w}({{\bf{f}}_{i}},{{\bf{f}}_{j}},{{\bf{S}}_{{{\cal C}_{p}}}}){\left({{y_{p,i}}-{y_{p,j}}}\right)^{2}},$ (8)
where ${\bf{f}}_{i}$ and ${\bf{f}}_{j}$ are feature vectors of node $i$ and node $j$, respectively, and weight $\textrm{w}({{\bf{f}}_{i}},{{\bf{f}}_{j}},{{\bf{S}}_{{{\cal C}_{p}}}})$ is introduced to measure a conformity between node $i$ and node $j$ in the feature space and similarity-based distribution. Intuitively, the pairwise potential encourages that as a higher weight $\textrm{w}({{\bf{f}}_{i}},{{\bf{f}}_{j}},{{\bf{S}}_{{{\cal C}_{p}}}})$ of node $i$ and node $j$, the closer refined similarity values between $y_{p,i}$ and $y_{p,j}$. These potentials are designed to reach a consensus from all the nodes on a clique for refining the initial similarities.
While we perform C-CRF on cliques to be robust against noisy edges, we further found that, in challenging datasets, a high number of noisy edges can exist even in the clique, deteriorating the similarity refinement process through C-CRFs. As a result, we design the weight function to be robust even in this extreme case by utilizing a statistical distance considering a similarity-based distribution in the clique.
More specifically, our probability mass function (PMF) in terms of a node $i$ is defined by $l_{2}$-normalization within a clique of a pivot node $p$, followed by a softmax as:
${Q_{i}}\left(k\right)=\frac{{\exp\left({{{\hat{s}}_{i,k}}}\right)}}{{\sum\nolimits_{l\in{\cal C}_{p}}{\exp\left({{{\hat{s}}_{i,l}}}\right)}}},\quad i,k\in{\cal C}_{p}$ (9)
where ${\hat{s}}_{i,k}$ is a $l_{2}$-normalized $s_{i,k}$ in terms of $k$.
We assume that two nodes have statistically similar distributions, if those two nodes are close to each other based on this PMF. In other words, PMF of a node is designed to serve as a descriptor of the node and, in this sense, we call it Similarity-Based Distribution (SBD). We then employ Jeffreys Divergence (J-Divergence) [22] symmetrizing Kullback-Leibler divergence for calculating the statistical distance between two distributions, as follows:
${D_{J}}\left({{Q_{i}}\parallel{Q_{j}}}\right)=\frac{1}{2}\left({{D_{KL}}\left({{Q_{i}}\parallel{Q_{j}}}\right)+{D_{KL}}\left({{Q_{j}}\parallel{Q_{i}}}\right)}\right),$ (10)
where each of $Q_{i},Q_{j}$ is a vector that has $L$ bins for all the other nodes in ${\cal C}_{p}$.
Finally, our weight function can be characterized with a Gaussian kernel of the Euclidean distance between two feature vectors ${\bf{f}}_{i}$ and ${\bf{f}}_{j}$, and our statistical distance:
$\textrm{w}({{\bf{f}}_{i}},{{\bf{f}}_{j}},{{\bf{S}}_{{C_{p}}}})=\exp\left\{{-\frac{{\left\|{{{\bf{f}}_{i}}-{{\bf{f}}_{j}}}\right\|_{2}^{2}}}{{2\sigma_{d}^{2}}}-\frac{{{D_{J}}\left({{Q_{i}}\parallel{Q_{j}}}\right)}^{2}}{{2\sigma_{r}^{2}}}}\right\},$ (11)
where hyper parameters, $\sigma_{d}$ and $\sigma_{r}$, adjust the degree of nearness. In the Gaussian kernel of our weight function, we call the first term as Euclidean Distance (ED) and the second term as Statistical Distance (SD). In an extreme case of calculating a weight between a node and its hard negative, our SD can identify the hard negative better than ED. This is because SD sees all the other nodes in the clique, some of which are properly represented in the feature space, even while the hard negative is not.
Multivariate Gaussian form. For ease of explanation, we now use a matrix notation within a clique of a specified size $L$ such as a weight matrix, ${\bf{W}}_{p}\in\mathbb{R}^{L\times L}$, whose element $w_{i,j}$ corresponds to $\textrm{w}({{\bf{f}}_{i}},{{\bf{f}}_{j}},{{\bf{S}}_{{{\cal C}_{p}}}})$. From now on, we also treat ${\bf{S}}_{{{\cal C}_{p}}}$ as a matrix that is composed of elements $s_{i,j}$ for the sake of simple explanation. A variant [23] of the C-CRF method simply represents Eq. 6 into a multivariate Gaussian form, as follows:
$\frac{1}{{{{\left({2\pi}\right)}^{\frac{L}{2}}}{{\left|\Sigma_{p}\right|}^{\frac{1}{2}}}}}\exp\left({-\frac{1}{2}{{\left({{{\bf{y}}_{p}}-{\mu_{p}}}\right)}^{T}}{\Sigma_{p}^{-1}}\left({{{\bf{y}}_{p}}-{\mu_{p}}}\right)}\right).$ (12)
The covariance matrix $\Sigma_{i}$ and mean $\mu_{i}$ is then defined as:
$\begin{array}[]{c}{\Sigma_{p}^{-1}}=2\left({\alpha{\bf{I}}+\beta{\bf{D}}_{p}-\beta{\bf{W}}_{p}}\right),\\ {\mu_{p}}=\Sigma{\bf{b}}_{p},\end{array}$ (13)
where ${\bf{b}}_{p}=2\alpha{\bf{s}}_{p}$, ${{\bf{s}}_{p}}={\left[{{s_{p,1}},...,{s_{p,L}}}\right]^{T}}$, and $D_{p}$ is the degree matrix of ${\bf{W}}_{p}$.
Inference. We can simply find ${\bf{y}}_{p}^{*}$ that maximizes the conditional probability ${\rm{P}_{p}}\left({{{\bf{y}}_{p}}|{{\bf{S}}_{{{\cal C}_{p}}}}}\right)$ by following the Gaussian property:
${\bf{y}}_{p}^{*}=\mathop{\arg\max}\limits_{{{\bf{y}}_{p}}}{{\rm{P}}_{p}}\left({{{\bf{y}}_{p}}|{{\bf{S}}_{{C_{p}}}}}\right)={\mu_{p}}.$ (14)
We then repeat to calculate the solution for the whole database images by setting each database image to a pivot image $I_{p}$. We can also employ a conjugate gradient (CG) method [24] to calculate the approximated solution of the inverse problem for computational efficiency. A constraint for CG is satisfied because the covariance matrix is symmetric and positive semi-definite.
Affinity matrix of denoised similarities. Existing diffusion methods [6, 12] employed the reciprocity check over $k$-NN lists for constructing an affinity matrix due to noises and outliers. On the other hand, our denoised similarities of $\bf{y}^{*}$ do not have much of noises and outliers. Thus we can instead use a simple symmetric affinity from the NN graph of our denoised similarities as $a_{pi}=\left(s_{k}(\mathbf{x}_{p}|\mathbf{x}_{i})+s_{k}(\mathbf{x}_{i}|\mathbf{x}_{p})\right)/2$. Afterward, we follow steps of each of diffusion methods with the denoised, symmetric affinity matrix to perform image retrieval.
3.2 Offline complexity
Thanks to the strategic use of cliques, we achieve significant computational efficiency in our C-CRF based denoising process. We outline the computational complexity by detailing each component involved in the process. For the inference stage of a pivot point, we perform three kinds of operations, such as $k$-NN for finding its clique, computing the weight matrix $\bf{W}$ within the clique and calculating the solution for similarity inference. The optimal $k$-NN search takes $\mathcal{O}(k\log N)$ with a whole database size $N$, and it can be accelerated by using some approximated methods [25] if $N$ is sufficiently large. Computing the weight matrix costs $\mathcal{O}(L^{3})$ with the specified clique size $L$ since J-Divergence (Eq. 10) takes $\mathcal{O}(L^{3})$ and lower order terms, such as Euclidean distance, $l_{2}$-normalization and softmax, are disregarded. For similarity inference, it takes $\mathcal{O}(L^{2}t)$ with a specified number of iterations, $t$, for the conjugate gradient method [24]. Note that complexities of these operations are bounded to the fixed clique size $L\ll N$ and the whole complexity for one pivot then becomes a constant time. We perform the same C-CRF refinement for each database image, but this can be parallelized and accelerated by using GPUs and CPUs.
4 Experiments
In this section, we explore the application of our denoising approach to two distinct but related tasks: landmark retrieval and person re-identification (re-ID). Both tasks benefit from robust image retrieval techniques, making them ideal for demonstrating the versatility of our method.
4.1 Experimental setup
Datasets and features. We conduct our experiments using challenging datasets: $\mathcal{R}$Oxford and $\mathcal{R}$Paris [26] for landmark retrieval, and DukeMTMC-reID [27] for re-ID. $\mathcal{R}$Oxford and $\mathcal{R}$Paris have three kinds of evaluation protocols, i.e., Easy, Medium, and Hard. Specifically, $\mathcal{R}$Oxford contains $4,993$ images and $70$ queries that represent particular Oxford landmarks and $\mathcal{R}$Paris consists of $6,322$ images and $70$ queries related to particular Paris landmarks. DukeMTMC-reID has 16,522 images with 702 identities for training, 2,228 images with 702 ids for query and 17,661 images for the gallery.
We employ GeM networks [28] of ResNet and VGG versions for landmark retrieval, which are seminal and easy to reproduce in a publicly available author’s implementation [^1] https://github.com/filipradenovic/cnnimageretrieval-pytorch. For re-ID, we leverage ResNet-50 [29], which is publicly available in a repository [^2] https://github.com/layumi/Person_reID_baseline_pytorch, for the baseline network.
Parameter setup. We simply fix C-CRF parameters $\alpha$, $\beta$ to $1$, $0.1$, respectively, for all the following experiments. Parameters $\sigma_{d}$, $\sigma_{r}$ of the Gaussian kernel are set to $0.8$, $2\times 10^{-4}$ for VGG and $0.9$, $3.5\times 10^{-4}$ for ResNet. We empirically choose clique sizes of $1,000$, $500$ and $150$ for $\mathcal{R}$Oxford, $\mathcal{R}$Paris and DukeMTMC-reID datasets, respectively due to the diverse distributions over different datasets. For other baseline methods, we mainly follow the same parameters of each method.
4.2 Results
Table 1: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Oxford. ResNet and VGG for extracting global features are fine-tuned networks. Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
Method Easy Medium Hard
ResNet NN-Search 84.2 65.4 40.1
NNS + AQE 81.9 67.1 42.7
EGT 81.8 65.4 42.5
+ Ours 85.5(+3.7) 73.2(+7.8) 50.8(+8.3)
Online diffusion 84.4 67.0 37.9
+ Ours 91.5(+7.1) 73.7(+6.3) 45.3(+7.4)
Offline adiffusion 88.2 69.9 41.1
+ Ours 92.4(+4.2) 76.1(+6.2) 50.3(+9.2)
VGG NN-Search 79.4 60.9 32.
NNS + AQE 86.3 69.1 41.1
EGT 88.3 71.6 44.9
+ Ours 87.5(-0.8) 72.4(+0.8) 46.7(+1.8)
Online diffusion 83.2 67.4 39.9
+ Ours 88.4(+5.2) 71.4(+4.0) 42.6(+2.7)
Offline diffusion 87.2 70.4 42.0
+ Ours 90.0(+2.8) 73.1(+2.7) 45.0(+3.0)
Table 2: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Paris. ResNet and VGG for extracting global features are fine-tuned networks. Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
Method Easy Medium Hard
ResNet NN-Search 91.6 76.7 55.2
NNS + AQE 93.6 82.3 63.9
EGT 92.8 82.7 68.6
+ Ours 92.7(-0.1) 83.3(+0.6) 69.6(+1.0)
Online diffusion 93.7 88.5 78.3
+ Ours 95.0(+2.3) 91.0(+2.5) 80.8(+2.5)
Offline adiffusion 94.2 87.9 77.6
+ Ours 95.1(+0.9) 89.3(+1.4) 78.7(+1.1)
VGG NN-Search 86.8 69.3 44.2
NNS + AQE 91.0 75.4 52.7
EGT 92.3 81.2 64.7
+ Ours 92.5(+0.2) 83.1(+1.9) 69.3(+4.6)
Online diffusion 92.7 85.9 74.2
+ Ours 94.7(+2.0) 89.1(+3.2) 78.2(+4.0)
Offline diffusion 92.2 84.1 72.3
+ Ours 93.3(+1.1) 85.4(+1.3) 73.2(+0.9)
Table 3: Performance Comparisons against with other re-ranking methods on re-ID setting of DukeMTMC-reID. Rank 1 means a cumulated accuracy at rank-1. Improvements without online-time cost are indicated in red.
Method Rank 1 mAP
NN-Search 64.6 43.6
k-reciprocal 69.3 60.7
EGT 64.5 47.5
+ Ours 64.6(+0.1) 48.2(+0.7)
Online diffusion 68.9 54.9
+ Ours 70.7(+1.8) 62.4(+7.5)
Offline diffusion 68.6 55.0
+ Ours 70.8(+2.2) 62.2(+7.2)
Complementarity Analysis. For validating the denoising quality, we exploit three re-ranking methods, i.e., online diffusion [6], offline diffusion [12], and Explore-Exploit Graph Traversal (EGT) [10]. Tables 1, 2, 3 show performance changes when our method is integrated with other re-ranking methods. Other tested method include Average Query Expansion (AQE) [4] on $R$Oxford and $R$Paris datasets, and k-reciprocal [30] for re-ID setting. In the comprehensive settings analyzed, our method predominantly demonstrated enhancements, with degradation occurring in only two cases, highlighting the high quality of our denoising approach. Overall, the average improvements with our method are 4.6 mAP for $\mathcal{R}$Oxford, 1.7 mAP for $\mathcal{R}$Paris and 3.3 mAP for DukeMTMC-reID. These results are mainly acquired thanks to our C-CRF based denoising.
Table 4: Ablation study of each term constituting our weight function with the Medium protocol, which can represent an overall quality, of each dataset. “w/ ED” and “w/ SD” represent to use the Euclidean distance of CNN-based features and the statistical distance of our similarity-based descriptors, respectively within our C-CRF. Baseline (top row in each feature category w/o any marks) indicates to use only diffusion without C-CRF. The best results are highlighted in bold.
w/ ED w/ SD Offline diffusion Online diffusion
$\mathcal{R}$Oxford $\mathcal{R}$Paris $\mathcal{R}$Oxford $\mathcal{R}$Paris
ResNet 69.9 87.9 67.0 88.5
75.5 89.2 71.4 89.8
75.9 89.3 73.0 91.0
76.1 89.3 73.7 91.0
VGG 70.4 84.1 67.4 85.9
70.9 84.9 67.8 86.2
72.7 85.3 71.0 88.9
73.1 85.4 71.4 89.1
Ablation study. We conduct ablation experiments for each term of the weight function of the C-CRF on both of online diffusion [6] and offline diffusion [12]. The experiments are conducted with the Medium protocol of $R$Oxford and $R$Paris, since the Medium protocol consisting of both of Easy and Hard images that can represent an overall quality of retrieval results. Table 4 displays performances with different weight terms on each network for feature extraction. As shown in the ablation study, we see that using both terms of the Euclidean Distance (ED) of CNN-based features and the Statistical Distance (SD) of similarity-based distributions for the weight function yields the best performance. We then use our C-CRF with ED and SD in the rest of experiments.
In Figure 3, we additionally experiment with measuring performance effect on varying $k$ parameters since a sparsity of the $k$-NN graph is determined from the $k$, and the sparsity affects the online time complexity; specifically, high density gives rise to high diffusion complexity. In the same vein, our resulting graph outperforms the reciprocal-NN graph with a more significant gap for smaller k (than 50), and this means we can reduce the online time complexity with less degradation of the retrieval quality. Furthermore, diffusion-based methods are known to be sensitive to the choice of parameters [31]. Our approach helps to mitigate the sensitivity to the important parameter $k$.
In Figure 4, the mAP performances are shown with varying clique sizes. The ablation study analyzes the performance according to the clique size, demonstrating robust results over a wide range of clique sizes compared to the baseline. We choose appropriate clique sizes based on this study, even though our method is robust to varying clique sizes in a wide range when compared to baseline.
(a) $R$Oxford-ResNet
(b) $R$Paris-ResNet
Figure 3: mAP graphs on varying $k$. The mAP results are measured on the Medium protocol that reflects the overall performance.
(a) $R$Oxford-ResNet
(b) $R$Paris-ResNet
Figure 4: mAP graphs on varying sizes of the clique. The mAP results are measured on the Medium protocol.
(a) Pivot image
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
4.3 Qualitative results
[…]
(a) Pivot image
0ptRefer to caption
(b) G-ED:0.36
G-SD:0.65
0ptRefer to caption
(c) G-ED:0.34
G-SD:0.68
0ptRefer to caption
(d) G-ED:0.33
G-SD:0.45
0ptRefer to caption
(e) G-ED:0.37
G-SD:0.35
0ptRefer to caption
(f) G-ED:0.35
G-SD:0.37
0ptRefer to caption
(g) G-ED:0.34
G-SD:0.27
Figure 6: Weights (conformities) used for C-CRF refinement between a pivot image and each of its positive and negative images. Red and blue bars denote negative and positive images of the pivot, respectively. Each of G-ED and G-SD is a weight through the Gaussian kernel (Eq. 11) computed by each distance. The precise weights can be helpful while attaining a right consensus through C-CRF. A higher value means more positive image from the pivot image.
5 Conclusion
In this study, we introduced a C-CRF based refinement method for denoising NN graph, which is crucial for visual re-ranking. By leveraging cliques, this method enhances computational efficiency while ensuring robust distance measurements through statistical distance metrics for C-CRF. Our approach significantly improves performance in landmark retrieval and person re-identification tasks, demonstrating its effectiveness as a pre-processing tool for enhancing visual re-ranking accuracy.

### v2

@@
-Denoising Nearest Neighbor Graph via Continuous CRF for Visual Re-ranking without Fine-tuning
+Enhancing Visual Re-ranking through Denoising Nearest Neighbor Graph via Continuous CRF
 Abstract
-Visual re-ranking using Nearest Neighbor graph (NN graph) has been adapted to yield high retrieval accuracy, since it is beneficial to exploring an high-dimensional manifold and applicable without additional fine-tuning.
-The quality of visual re-ranking using NN graph, however, is limited to that of connectivity, i.e., edges of the NN graph.
-Some edges can be misconnected with negative images.
-This is known as a noisy edge problem, resulting in a degradation of the retrieval quality.
-To address this, we propose a complementary denoising method based on Continuous Conditional Random Field (C-CRF) that uses a statistical distance of our similarity-based distribution.
-This method employs the concept of cliques to make the process computationally feasible.
-We demonstrate the complementarity of our method through its application to three visual re-ranking methods, observing quality boosts in landmark retrieval and person re-identification (re-ID).
+Nearest neighbor (NN) graph based visual re-ranking has emerged as a powerful approach for improving retrieval accuracy, offering the advantages of effectively exploring high-dimensional manifolds without requiring additional fine-tuning.
+However, the effectiveness of NN graph-based re-ranking is fundamentally constrained by the quality of its edge connectivity, as incorrect connections between dissimilar (negative) images frequently occur.
+This is known as a noisy edge problem, which hinders the re-ranking performance of existing techniques and limits their potential.
+To remedy this issue, we propose a complementary denoising method based on Continuous Conditional Random Fields (C-CRF) that leverages statistical distances derived from similarity-based distributions.
+As a pre-processing step for enhancing NN graph-based retrieval, our approach constructs fully connected cliques around each anchor image and employs a novel statistical distance metric to robustly alleviate noisy edges before re-ranking while achieving efficient processing through offline computation.
+Extensive experimental results demonstrate that our method consistently improves three different NN graph-based re-ranking approaches, yielding significant gains in retrieval accuracy.
+Index Terms—  Visual re-ranking, Visual retrieval, Nearest neighbor search
 1 Introduction
 (a) Initial NN graph
-(b) NN graph after reciprocity check
-(c) Diffusion-based retrieval of (b)
-(d) NN graph denoised by ours
-(e) Diffusion-based retrieval of (d)
-Figure 1: This figure shows NN graph (affinity) of toy data on manifold.
+(b) NN graph after
+reciprocity check
+(c) Diffusion-based
+retrieval of (b)
+(d) NN graph denoised by
+our method
+(e) Diffusion-based
+retrieval of (d)
+Fig.
+1: This figure shows NN graph (affinity) of toy data on manifold.
 The blue and green colors of nodes indicate different manifolds, while a yellow node and red nodes represent a query point and the retrieval results from a visual re- ranking method of diffusion.
@@
 Results (e) of our approach (d) are on the correct manifold.
-Best viewed with zoom-in.
-Nearest neighbor search approaches  [1, 2, 3] have used various visual re-ranking techniques  [4, 5, 6] for further improvement of retrieval accuracy.
-Among them, average query expansion (AQE)  [4] was considered to explore an image manifold on the descriptor space in a simple manner by using a new query averaged over nearest neighbor features at a query time.
-However, AQE is limited to only information between query and database images.
-Contrary to AQE, NN graph based retrieval methods, e.g., diffusion process [7, 8, 6, 9] and graph traversal [10, 11], fully makes use of an affinity matrix between database images as well as information between the query and database images.
-Diffusion methods, revisited from Donoser et al.
-[8], constructed the affinity matrix, while considering the query as a part of database images, and calculated solutions of diffusion methods in an iterative manner.
-However, regarding the query as one of database images requires huge computational costs in online time, because of creating a new affinity matrix whenever new queries are given.
-Iscen et al.
-[6] proposed a method that does not have to recreate the matrix even if new queries are given.
-Yang et al.
-[12] also made the Iscen’s method more efficient by shifting the diffusion step from online computing to offline one.
-Recently, Ouyang et al.
-[13] utilizes self-attention modules to aggregate similarities for improving visual re-ranking outcomes, while requiring additional fine-tuning.
-Shao et al.
-[14] highlight visual re-ranking from only top-M results using rich global features to enhance only retrieval precision, not recall.
-The reliance of this method on the initial ranked list may reduce its robustness across varied or broader retrieval contexts.
-Our work assumes conditions without fine-tuning and without rich global features, performing re-ranking across the entire database [15].
-Yang et al.
-[16] tackled an issue where NN graph based re-ranking is quite sensitive to noisy data, since all edges, even those containing noise, are used to represent the affinity matrix.
-Thus, a locally constrained affinity construction was proposed based on edges of satisfying $k$ nearest neighbors ($k$-NN) constraints of images for relaxing noisy edges of negative images.
-A seminal approach [6] suggested an extension by filtering out noisy edges even from the locally constrained affinity by using a reciprocity check of outgoing and incoming edges of the locally constrained affinity (NN graph).
-In this paper, we found that this reciprocity check with NN graph is insufficient to address the noisy edge problem, since the reciprocity check considers only reciprocal relationships between two edges.
-Since the reciprocity check and $k$-NN constraints do not fully exploit all information from each other, we address the problem with full consideration of relationships between all edges in a clique where we can conservatively reach a consensus (Fig.
-1).
-To realize this, we efficiently employ a fully connected C-CRF, when denoising the noisy NN graph in offline time.
-Conditional Random Field (CRF) has been extensively studied [17], and Continuous CRF (C-CRF) [18] was proposed for regression tasks such as depth estimation, image denoising and metric learning for person re-ID [19, 20, 21].
-Departing to these applications, our approach targets to refine similarities between each of database images that are robust to noisy edges for improving a diffusion performance in offline time.
-Main contributions are summarized as follows:
-  • We introduce a C-CRF based refinement method for noisy NN graphs that serves as a pre-processing module to enhance visual re-ranking.
-  • Our approach utilizes a statistical distance metric to ensure robust distance measurement and employs cliques for cost-effective denoising.
-  • We show that our method significantly enhances existing visual re-ranking methods in landmark retrieval and person re-identification, confirming its role as a complementary tool.
-2 Visual re-ranking with NN graph
+Visual retrieval systems typically perform nearest neighbor search with visual descriptors; however, they face fundamental challenges when handling non-linear manifold structures in high-dimensional feature spaces [1, 2].
+To address this issue, re-ranking process has emerged as a crucial component in visual retrieval systems, refining initial search results to improve retrieval accuracy.
+By leveraging various contextual information and relationships between images, re-ranking methods have consistently demonstrated their ability to enhance the quality of retrieval results beyond the initial ranking [3, 4, 5].
+Visual re-ranking techniques have evolved along several distinct research directions.
+Geometric verification approaches [4, 6, 7] leverage local descriptors to verify spatial consistency between image pairs, effectively reducing false matches but incurring significant computational overhead.
+In contrast, query expansion methods [8, 5, 9] utilize global descriptors to explore image manifolds, aggregating nearest neighbor features to refine the original query descriptor for subsequent retrieval processes.
+Learning-based approaches [10, 11] employ dedicated re-ranking models with neural networks, however, these methods typically require additional fine-tuning or training data.
+Meanwhile, nearest neighbor graph-based methods [12, 13] have gained prominence for their ability to explore high-dimensional manifolds without requiring additional training.
+These approaches, particularly diffusion processes [14, 3] and graph traversal methods [15, 16], achieve superior performance by effectively utilizing comprehensive affinity relationships within the image database.
+However, the effectiveness of nearest neighbor graph-based re-ranking is fundamentally constrained by the quality of its edge connectivity in the graph structure.
+Specifically, incorrect connections between dissimilar (negative) images frequently occur in the initial nearest neighbor graph, known as the noisy edge problem.
+Such noisy edges significantly hinder re-ranking performance by introducing false connectivity paths that propagate incorrect similarity information throughout the manifold structure.
+While previous works have attempted to address this issue through locally constrained affinity construction [17] and reciprocity checks [14], these solutions are limited as they only consider pairwise relationships between edges.
+The reciprocity check, in particular, fails to fully exploit collective information across the graph, as it only considers mutual neighbor relationships while ignoring broader neighborhood structures, leading to suboptimal performance in complex manifold scenarios.
+To overcome these limitations, we propose a complementary denoising approach based on Continuous Conditional Random Fields (C-CRF) [18] that leverages statistical distances derived from similarity-based distributions.
+While CRF has been extensively studied for error reduction in various computer vision tasks such as segmentation and depth estimation [19, 20, 21], its potential for addressing the noisy edge problem in visual retrieval remains unexplored.
+As a pre-processing step for enhancing NN graph-based retrieval, our method constructs fully connected cliques around each anchor image and considers relationships among all edges within these cliques to reach a more robust consensus, effectively addressing the limitations of pairwise verification approaches.
+Unlike previous applications of CRF, our approach uniquely focuses on refining pairwise similarities between database images through efficient offline processing, making them robust against noisy edges and thereby improving re-ranking performance.
+Extensive experimental results demonstrate that our method consistently enhances three different nearest neighbor graph-based re-ranking approaches, yielding significant improvements in retrieval accuracy.
+In summary, our contributions are as follows:
+  • We propose a complementary denoising method based on C-CRF that enhances NN graph-based re-ranking without requiring additional networks or fine-tuning.
+  • We design a statistical distance metric derived from similarity-based distributions and employ fully connected cliques for efficient offline graph denoising.
+  • Experimental results demonstrate that our method consistently improves three different NN graph-based re-ranking approaches, confirming its effectiveness as an easily integrable complementary tool.
+2 Preliminaries
 In this section, we summarize the diffusion process to show how the NN-graph can be applied for visual re-ranking.
@@
 2).
-Figure 2: An example of clique-based denoising; $k=7$ and $L=10$.
+Fig.
+2: An example of clique-based denoising; $k=7$ and $L=10$.
 (a) shows initial edges of a yellow pivot node, and red edges are noisy edges.
@@
 Our goal is to refine initial similarities of NN graph that are commonly set by top-$k$ similarities of each database image.
-For computational efficiency and effective noise handling, we utilize the concept of cliques and apply C-CRF into each of cliques.
+For computational efficiency and effective noise handling, we utilize the concept of cliques rather than processing the entire graph.
+Each clique is then treated with C-CRF to optimize performance.
 Specifically, we choose a clique, a complete subgraph, where we can have ample information such that we can identify noisy edges.
-On top of that, using the clique has computational benefits in running CRFs (Sec.
-3.2).
 We suppose that the whole database size is $N$ and a clique of a pivot image $\left\{{{I_{p}}}\right\}_{p=1}^{N}$ is ${{\cal C}_{p}}=\left({{{\cal N}_{p}},{{\cal E}_{p}}}\right)$, where a set of nodes (images) ${\cal N}_{p}$ is composed of $L$ nearest neighbors from the pivot image $I_{p}$ by a specified clique size $L$, and ${\cal E}_{p}$ is a set of edges that makes the clique ${\cal C}_{p}$, i.e., a complete subgraph.
@@
 On the other hand, our denoised similarities of $\bf{y}^{*}$ do not have much of noises and outliers.
-Thus we can instead use a simple symmetric affinity from the NN graph of our denoised similarities as $a_{pi}=\left(s_{k}(\mathbf{x}_{p}|\mathbf{x}_{i})+s_{k}(\mathbf{x}_{i}|\mathbf{x}_{p})\right)/2$.
-Afterward, we follow steps of each of diffusion methods with the denoised, symmetric affinity matrix to perform image retrieval.
-3.2 Offline complexity
-Thanks to the strategic use of cliques, we achieve significant computational efficiency in our C-CRF based denoising process.
-We outline the computational complexity by detailing each component involved in the process.
-For the inference stage of a pivot point, we perform three kinds of operations, such as $k$-NN for finding its clique, computing the weight matrix $\bf{W}$ within the clique and calculating the solution for similarity inference.
-The optimal $k$-NN search takes $\mathcal{O}(k\log N)$ with a whole database size $N$, and it can be accelerated by using some approximated methods [25] if $N$ is sufficiently large.
-Computing the weight matrix costs $\mathcal{O}(L^{3})$ with the specified clique size $L$ since J-Divergence (Eq.
-10) takes $\mathcal{O}(L^{3})$ and lower order terms, such as Euclidean distance, $l_{2}$-normalization and softmax, are disregarded.
-For similarity inference, it takes $\mathcal{O}(L^{2}t)$ with a specified number of iterations, $t$, for the conjugate gradient method [24].
-Note that complexities of these operations are bounded to the fixed clique size $L\ll N$ and the whole complexity for one pivot then becomes a constant time.
-We perform the same C-CRF refinement for each database image, but this can be parallelized and accelerated by using GPUs and CPUs.
+Thus we can instead use a simple symmetric affinity from the NN graph of our denoised similarities as $a_{pi}=\left({\bf{y}}_{p,i}^{*}+{\bf{y}}_{i,p}^{*}\right)/2$.
+Afterward, we follow steps of each of diffusion methods with the denoised, symmetric affinity matrix to perform visual re-ranking.
 4 Experiments
-In this section, we explore the application of our denoising approach to two distinct but related tasks: landmark retrieval and person re-identification (re-ID).
-Both tasks benefit from robust image retrieval techniques, making them ideal for demonstrating the versatility of our method.
+In this section, we explore the application of our denoising approach to landmark retrieval with three visual re-ranking methods.
 4.1 Experimental setup
 Datasets and features.
-We conduct our experiments using challenging datasets: $\mathcal{R}$Oxford and $\mathcal{R}$Paris [26] for landmark retrieval, and DukeMTMC-reID [27] for re-ID.
+We conduct our experiments using challenging datasets: $\mathcal{R}$Oxford and $\mathcal{R}$Paris [26] for landmark retrieval.
 $\mathcal{R}$Oxford and $\mathcal{R}$Paris have three kinds of evaluation protocols, i.e., Easy, Medium, and Hard.
 Specifically, $\mathcal{R}$Oxford contains $4,993$ images and $70$ queries that represent particular Oxford landmarks and $\mathcal{R}$Paris consists of $6,322$ images and $70$ queries related to particular Paris landmarks.
-DukeMTMC-reID has 16,522 images with 702 identities for training, 2,228 images with 702 ids for query and 17,661 images for the gallery.
-We employ GeM networks [28] of ResNet and VGG versions for landmark retrieval, which are seminal and easy to reproduce in a publicly available author’s implementation [^1] https://github.com/filipradenovic/cnnimageretrieval-pytorch.
-For re-ID, we leverage ResNet-50 [29], which is publicly available in a repository [^2] https://github.com/layumi/Person_reID_baseline_pytorch, for the baseline network.
+We employ GeM networks [9] of ResNet and VGG versions for landmark retrieval, which are easy to reproduce in a publicly available author’s implementation [^1] https://github.com/filipradenovic/cnnimageretrieval-pytorch.
 Parameter setup.
@@
 Parameters $\sigma_{d}$, $\sigma_{r}$ of the Gaussian kernel are set to $0.8$, $2\times 10^{-4}$ for VGG and $0.9$, $3.5\times 10^{-4}$ for ResNet.
-We empirically choose clique sizes of $1,000$, $500$ and $150$ for $\mathcal{R}$Oxford, $\mathcal{R}$Paris and DukeMTMC-reID datasets, respectively due to the diverse distributions over different datasets.
+We empirically choose clique sizes of $1,000$ and $500$ for $\mathcal{R}$Oxford and $\mathcal{R}$Paris datasets, respectively.
 For other baseline methods, we mainly follow the same parameters of each method.
 4.2 Results
-Table 1: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Oxford.
+Table 1: mean Average Precision (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Oxford.
 ResNet and VGG for extracting global features are fine-tuned networks.
@@
 + Ours 90.0(+2.8) 73.1(+2.7) 45.0(+3.0)
-Table 2: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Paris.
-ResNet and VGG for extracting global features are fine-tuned networks.
-Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
+Table 2: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Paris. The layout is the same as in Table 1.
 Method Easy Medium Hard
@@
 + Ours 93.3(+1.1) 85.4(+1.3) 73.2(+0.9)
-Table 3: Performance Comparisons against with other re-ranking methods on re-ID setting of DukeMTMC-reID.
-Rank 1 means a cumulated accuracy at rank-1.
-Improvements without online-time cost are indicated in red.
-Method Rank 1 mAP
-NN-Search 64.6 43.6
-k-reciprocal 69.3 60.7
-EGT 64.5 47.5
-+ Ours 64.6(+0.1) 48.2(+0.7)
-Online diffusion 68.9 54.9
-+ Ours 70.7(+1.8) 62.4(+7.5)
-Offline diffusion 68.6 55.0
-+ Ours 70.8(+2.2) 62.2(+7.2)
 Complementarity Analysis.
 For validating the denoising quality, we exploit three re-ranking methods, i.e., online diffusion [14], offline diffusion [22], and Explore-Exploit Graph Traversal (EGT) [15].
-Tables 1, 2, 3 show performance changes when our method is integrated with other re-ranking methods.
-Other tested method include Average Query Expansion (AQE) [4] on $R$Oxford and $R$Paris datasets, and k-reciprocal [30] for re-ID setting.
-In the comprehensive settings analyzed, our method predominantly demonstrated enhancements, with degradation occurring in only two cases, highlighting the high quality of our denoising approach.
-Overall, the average improvements with our method are 4.6 mAP for $\mathcal{R}$Oxford, 1.7 mAP for $\mathcal{R}$Paris and 3.3 mAP for DukeMTMC-reID.
+Tables 1, 2 show performance changes when our method is integrated with other re-ranking methods.
+Other tested method include Average Query Expansion (AQE) [8] on $R$Oxford and $R$Paris datasets.
+Overall, the average improvements with our method are 4.6 mAP for $\mathcal{R}$Oxford, 1.7 mAP for $\mathcal{R}$Paris.
 These results are mainly acquired thanks to our C-CRF based denoising.
-Table 4: Ablation study of each term constituting our weight function with the Medium protocol, which can represent an overall quality, of each dataset.
-“w/ ED” and “w/ SD” represent to use the Euclidean distance of CNN-based features and the statistical distance of our similarity-based descriptors, respectively within our C-CRF.
-Baseline (top row in each feature category w/o any marks) indicates to use only diffusion without C-CRF.
+Table 3: Ablation study of each term constituting our weight function.
+“w/ ED” and “w/ SD” represent to use the Euclidean distance of CNN-based features and the statistical distance of our similarity-based descriptors, respectively.
+Baseline indicates to use only diffusion without C-CRF.
 The best results are highlighted in bold.
-w/ ED w/ SD Offline diffusion Online diffusion
-$\mathcal{R}$Oxford $\mathcal{R}$Paris $\mathcal{R}$Oxford $\mathcal{R}$Paris
-ResNet 69.9 87.9 67.0 88.5
-75.5 89.2 71.4 89.8
-75.9 89.3 73.0 91.0
-76.1 89.3 73.7 91.0
-VGG 70.4 84.1 67.4 85.9
-70.9 84.9 67.8 86.2
-72.7 85.3 71.0 88.9
-73.1 85.4 71.4 89.1
+w/ ED w/ SD $\mathcal{R}$Oxford $\mathcal{R}$Paris
+ResNet 69.9 87.9
+75.5 89.2
+75.9 89.3
+76.1 89.3
+VGG 70.4 84.1
+70.9 84.9
+72.7 85.3
+73.1 85.4
 Ablation study.
-We conduct ablation experiments for each term of the weight function of the C-CRF on both of online diffusion [6] and offline diffusion [12].
+We conduct ablation experiments for each term of the weight function of the C-CRF on offline diffusion [22].
 The experiments are conducted with the Medium protocol of $R$Oxford and $R$Paris, since the Medium protocol consisting of both of Easy and Hard images that can represent an overall quality of retrieval results.
 Table 3 displays performances with different weight terms on each network for feature extraction.
-As shown in the ablation study, we see that using both terms of the Euclidean Distance (ED) of CNN-based features and the Statistical Distance (SD) of similarity-based distributions for the weight function yields the best performance.
-We then use our C-CRF with ED and SD in the rest of experiments.
+As shown in the ablation study, we see that using both terms of the Euclidean Distance (ED) of CNN features and the Statistical Distance (SD) of similarity distributions for the weight function yields the best performance.
 In Figure 3, we additionally experiment with measuring performance effect on varying $k$ parameters since a sparsity of the $k$-NN graph is determined from the $k$, and the sparsity affects the online time complexity; specifically, high density gives rise to high diffusion complexity.
@@
 (b) $R$Paris-ResNet
-Figure 3: mAP graphs on varying $k$. The mAP results are measured on the Medium protocol that reflects the overall performance.
+Fig. 3: mAP graphs on varying $k$. The mAP results are measured on the Medium protocol.
 (a) $R$Oxford-ResNet
 (b) $R$Paris-ResNet
-Figure 4: mAP graphs on varying sizes of the clique. The mAP results are measured on the Medium protocol.
-(a) Pivot image
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-0ptRefer to caption
-[…]
-[…]
-[…]
-4.3 Qualitative results
-[…]
-[…]
-(a) Pivot image
-0ptRefer to caption
-(b) G-ED:0.36
-G-SD:0.65
-0ptRefer to caption
-(c) G-ED:0.34
-G-SD:0.68
-0ptRefer to caption
-(d) G-ED:0.33
-G-SD:0.45
-0ptRefer to caption
-(e) G-ED:0.37
-G-SD:0.35
-0ptRefer to caption
-(f) G-ED:0.35
-G-SD:0.37
-0ptRefer to caption
-(g) G-ED:0.34
-G-SD:0.27
-Figure 6: Weights (conformities) used for C-CRF refinement between a pivot image and each of its positive and negative images.
-Red and blue bars denote negative and positive images of the pivot, respectively.
-Each of G-ED and G-SD is a weight through the Gaussian kernel (Eq.
-11) computed by each distance.
-The precise weights can be helpful while attaining a right consensus through C-CRF.
-A higher value means more positive image from the pivot image.
+Fig. 4: mAP graphs on varying sizes of the clique. The mAP results are measured on the Medium protocol.
 5 Conclusion
-In this study, we introduced a C-CRF based refinement method for denoising NN graph, which is crucial for visual re-ranking.
-By leveraging cliques, this method enhances computational efficiency while ensuring robust distance measurements through statistical distance metrics for C-CRF.
-Our approach significantly improves performance in landmark retrieval and person re-identification tasks, demonstrating its effectiveness as a pre-processing tool for enhancing visual re-ranking accuracy.
+We propose a C-CRF-based denoising approach to enhance NN graph-based re-ranking.
+Our method leverages fully connected cliques and a novel statistical distance metric to robustly alleviate noisy edges.
+Experimental results demonstrate consistent improvements over existing re-ranking approaches, confirming the effectiveness of our method as a complementary tool for enhancing visual re-ranking.
+Acknowledgements
+This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant (No.
+RS-2023-00237965, Recognition, Action and Interaction Algorithms for Open-world Robot Service) and the National Research Foundation of Korea (NRF) grant (No.
+RS-2023-00208506), both funded by the Korea government (MSIT).
+Prof.
+Sung-Eui Yoon is a corresponding author (e-mail: sungeui@kaist.edu).